Trico Company set the following standard unit costs for its single product.

Direct materials (30 Ibs. @ $5.10 per Ib.) $ 153.00
Direct labor (8 hrs. @ $14 per hr.) 112.00
Factory overhead—variable (8 hrs. @ $6 per hr.) 48.00
Factory overhead—fixed (8 hrs. @ $12 per hr.) 96.00
Total standard cost $ 409.00


The predetermined overhead rate is based on a planned operating volume of 80% of the productive capacity of 66,000 units per quarter. The following flexible budget information is available.

Operating Levels
70% 80% 90%
Production in units 46,200 52,800 59,400
Standard direct labor hours 369,600 422,400 475,200
Budgeted overhead
Fixed factory overhead $ 5,068,800 $ 5,068,800 $ 5,068,800
Variable factory overhead $ 2,217,600 $ 2,534,400 $ 2,851,200


During the current quarter, the company operated at 90% of capacity and produced 59,400 units of product; actual direct labor totaled 369,400 hours. Units produced were assigned the following standard costs.

Direct materials (1,782,000 Ibs. @ $5.10 per Ib.) $ 9,088,200
Direct labor (475,200 hrs. @ $14 per hr.) 6,652,800
Factory overhead (475,200 hrs. @ $18 per hr.) 8,553,600
Total standard cost $ 24,294,600


Actual costs incurred during the current quarter follow.

Direct materials (1,387,000 Ibs. @ $6.70 per lb.) $ 9,292,900
Direct labor (369,400 hrs. @ $11.60 per hr.) 4,285,040
Fixed factory overhead costs 3,196,500
Variable factory overhead costs 3,466,700
Total actual costs $ 20,241,140

Required:
1. Compute the direct materials cost variance, including its price and quantity variances.

AQ = Actual Quantity
SQ = Standard Quantity
AP = Actual Price
SP = Standard Price



2. Compute the direct labor cost variance, including its rate and efficiency variances.

AH = Actual Hours
SH = Standard Hours
AR = Actual Rate
SR = Standard Rate



3. Compute the overhead controllable and volume variances.
Answer #1

1) Direct materials cost variance = Std material cost - Actual material cost

= $9,088,200 - $9,292,900 = -$204,700 Unfavorable

Direct material price variance = (SP - AP)*AQ

= ($5.10 - $6.70)*1,387,000 = -$2,219,200 Unfavorable

Direct materials efficiency variance = (SQ - AQ)*SP

= (1,782,000 - 1,387,000)*$5.10 = $2,014,500 Favorable

2) Direct Labor Cost Variance = Std Labor Cost - Actual Labor Cost

= $6,652,800 - $4,285,040 = $2,367,760 Favorable

Direct Labor Rate Variance = (SR - AR)*AH

= ($14 - $11.60)*369,400 = $886,560 Favorable

Direct Labor Efficiency Variance = (SH - AH)*SR

= (475,200 - 369,400)*$14 = $1,481,200 Favorable

3) Variable OH Controllable Variance = (Std VOH rate - Actual VOH rate)*AH

Actual VOH Rate = (Actual Variable OH/Actual Hrs)

= $3,466,700/369,400 hrs = $9.38468 per hour

VOH Controllable Variance = ($6 - $9.38468)*369,400 = -$1,250,300 Unfavorable

VOH Volume Varaince = (SH - AH)*S VOH rate

= (475,200 - 369,400)*$6 = $634,800 Favorable

Fixed OH Controllable Variance = Budgeted OH - Actual Fixed OH

= $5,068,800 - $3,196,500 = $1,872,300 Favorable

Fixed OH Volume Variance = Applied OH for Std hrs - Budgeted OH

= (475,200*$12) - $5,068,800 = $633,600 Favorable